A particular species of plant grows 3.7 m in 14 days. What is it's growth rate in micrometers per seconds

1
Expert's answer
2022-01-30T13:53:07-0500

The growth rate is:


"r = \dfrac{3.7m}{14days} = \dfrac{3.7}{14}\dfrac{m}{days}\\"

In micrometers per seconds:


"r = \dfrac{3.7}{14}\cdot \dfrac{10^{6}\mu m}{86400s} \approx 3.1\mu m/s"

Answer. "3.1\mu m/s".